OpenStudy (anonymous):

write the equation -2x=-4y+12 in slope-intercept form.

OpenStudy (anonymous):

I got y=1/2x-3 is this correct?

OpenStudy (anonymous):

-2x = -4y + 12 4y = 2x + 12 y = 1/2 x + 3

OpenStudy (anonymous):

thank you sir